OpenStudy (anonymous):

I'm struggling with a simple decimal division for a speed distance time problem, speed = 6.3km/Hr Dist = 11.7km. No problem shifting the decimal point to make the whole numbers 63 and 117. 63 goes into 117, 1 remainder 54 and I'm stuck from here, please help this is driving me insane. I can see that 9 goes into 63, 7 times and 13 x 9 is 117 making 9 the common factor but I don't know how to get the answer down into 'hours minutes and seconds' - thanks to Unkle for getting me as far as he did

OpenStudy (nubeer):

speed = distance /time 6.3 = 11.7* 1/t \[63\div10 =117\div10 \times 1div t\] 10 will be cancelled out. t = 13/7 (9 is the factor)

OpenStudy (anonymous):

I've done this on the caculator and see the answer is 1.85, I can't seem to arrive at this without the calculator as I think the remainder is 54?

OpenStudy (unklerhaukus):

\[s = 6.3[\text{km/hr}]\qquad\qquad d=11.7 [\text {km}]\]\[s=\frac dt\] \[\qquad t=\frac ds\] \[\quad t =\frac{11.7 [\text {km}]}{6.3[\text{km/hr}]} \]\[\qquad=\frac{117}{63}\]\[\qquad=\frac{9\times 13}{9\times 7}\]\[\qquad=\frac{}{}\]

OpenStudy (anonymous):

9x13 = 117 ------ 9x7 = 63 ok?

OpenStudy (unklerhaukus):

\[\frac{117}{63}=\frac{63+54}{63}=1+\frac{54}{63}\]

OpenStudy (anonymous):

Ah I get it, 117 is the result of 13x9 but why did you break it down then to 9x7? And then why 63+54 ( is it becasuse this is the next closest step in the 9 times table?)

OpenStudy (unklerhaukus):

yeah use the nine times table

OpenStudy (unklerhaukus):

54 = 9 x ?

OpenStudy (anonymous):

6

OpenStudy (unklerhaukus):

yeah 9 x 6 =54 , 63= 9 x ?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

so the way to work it out is to find which times tablw the initial numbers fit into?

OpenStudy (unklerhaukus):

\[1+\frac{54}{63}=1+\frac{9\times6}{9\times7}=1+\frac{\cancel{9\times}6}{\cancel{9\times}7}=1+\frac{6}{7}\]

OpenStudy (unklerhaukus):

cancel the common factors

OpenStudy (anonymous):

ok I'm stuck with how to solve this portion, the answer is 1 remainder 1?

OpenStudy (unklerhaukus):

6/7 ~ 8.571. . ...

OpenStudy (unklerhaukus):

what form is your answer supposed to be /

OpenStudy (anonymous):

the answer is supposed to be in hours minutes seconds?

OpenStudy (unklerhaukus):

ohhhh,

OpenStudy (unklerhaukus):

did i guess right that speed was given in km/hr ?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

then we have 1 + 6/7 hours

OpenStudy (anonymous):

sorry should have made that clearer, thatnks for your time here

OpenStudy (unklerhaukus):

6/7 hours are 60 x 60/7 mins

OpenStudy (unklerhaukus):

360/ 7 mins

OpenStudy (unklerhaukus):

but we should factorise 360

OpenStudy (anonymous):

ok, I am totally lost, sorry

OpenStudy (anonymous):

apologies, I just lost it trying to get it into 'time'

OpenStudy (unklerhaukus):

its not a round number

OpenStudy (anonymous):

I know the answer is 1 hour 51 minutes 4 seconds but I have no idea how to get to that answer?